OpenStudy (timaashorty):

Can someone help me real quick? Given f(x)=x^2+7 and g(x)=x-4/x. Find (g°f)(-1) Hint: (g°f)(-4) = g(f(-4)

OpenStudy (timaashorty):

@SolomonZelman ?

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ f(x)=x^2+7 \LARGE\color{white}{ \rm │ }}\) \(\large\color{black}{ g(x)=\frac{x-4}{x} \LARGE\color{white}{ \rm │ }}\) \(\large\color{black}{ (gºf)(x)=\frac{(x^2+7)-4}{(x^2+7)} \LARGE\color{white}{ \rm │ }}\) \(\large\color{black}{ (gºf)(x)=\frac{x^2+3}{x^2+7} \LARGE\color{white}{ \rm │ }}\) \(\large\color{black}{ (gºf)(-1)=\frac{(-1)^2+3}{(-1)^2+7} \LARGE\color{white}{ \rm │ }}\) \(\large\color{black}{ (gºf)(-1)=\frac{1+3}{1+7} \LARGE\color{white}{ \rm │ }}\) \(\large\color{black}{ (gºf)(-1)=\frac{4}{8} \LARGE\color{white}{ \rm │ }}\) \(\large\color{black}{ (gºf)(-1)=\frac{1}{2} \LARGE\color{white}{ \rm │ }}\)

OpenStudy (solomonzelman):

IF you have any questions about the process , please ask.

OpenStudy (timaashorty):

Wow, love the way you gave me them step by step. helps a whole lot. thank you!